OpenStudy (ray10):

Find particular solution \[\large y''+y'+y=14+8x+11\cos(x)\]

OpenStudy (ray10):

I have tackled similar questions before, but this one is a bit of a challenge to me and was hoping for a in-depth insight :)

OpenStudy (unklerhaukus):

the first step is to find the corresponding homogenous solution with the auxiliary equation

OpenStudy (unklerhaukus):

\[m^2+m+1=0\]

OpenStudy (ray10):

@Mimi_x3 variation of parameters?

OpenStudy (ray10):

yes I had \[\large m^{2}+m+1=0\] as my characteristic equation @UnkleRhaukus as equation (1) I had \[\large 14+8x+11\cos(x)\] as equation (2) then I get lost

OpenStudy (unklerhaukus):

work out the two values of m using the quadratic formula

OpenStudy (ray10):

that comes out as \[\frac{-1\pm \sqrt3 i}{2}\]

OpenStudy (unklerhaukus):

yeah

OpenStudy (unklerhaukus):

you have complex solutions for m , so m=a±ib so the corresponding homogenous solution is \[y_c=e^{ax}\Big(A\cos(bx)+B\sin(bx)\Big)\]

OpenStudy (unklerhaukus):

What do you get for the homogenous solution @Ray10 ?

OpenStudy (ray10):

ooh okay, I haven't used this one before

OpenStudy (ray10):

so \[\large m = \frac{-1 \pm \sqrt3 i}{2}\]

OpenStudy (unklerhaukus):

yes so what is a, b?

OpenStudy (ray10):

a would be \[\large \frac{-1+\sqrt3 i}{2}\] ?? Or I am confused if not :S

OpenStudy (unklerhaukus):

um no, break the m up into a real term ± and imaginary term

OpenStudy (ray10):

oh now I get it!

OpenStudy (ray10):

a= \[\large \frac{-1}{2}\] and b= \[\large \frac { \pm \sqrt3 i}{2}\]

OpenStudy (unklerhaukus):

yeah thats pretty right, but the form was m=a±ib so the b term dosn't have the i in it

OpenStudy (unklerhaukus):

so now just sub in a=-1/2 and b=√3/2 into \[y_c=e^{ax}\Big(A\cos(bx)+B\sin(bx)\Big)\]

OpenStudy (ray10):

if it doesn't have an i in it does that make it: \[\large m= \frac{-1 \pm \sqrt3}{2}, \frac{-1}{2} \pm \frac{ \pm \sqrt3}{2}\] ? in the form of m=a \[a\pm ib\]

OpenStudy (ray10):

like where does the minus of the \[\pm\] go?

OpenStudy (unklerhaukus):

well you had m right, m does have an i in it, \[m=a\pm i b\] once we have this form we extract the real and plus/minus imaginary parts the formula for the complementary homogenous solution, only needs a, and b

OpenStudy (ray10):

\[y_{c}=e^{\frac{-1}{2}x}(Acos(\frac{\sqrt3}{2}x))+Bsin(\frac{\sqrt3}{2}x))\] ? :)

OpenStudy (ray10):

ah now that makes sense :)

OpenStudy (unklerhaukus):

YES! that's it

OpenStudy (ray10):

:D

OpenStudy (unklerhaukus):

So you have the complementary homogenous solution

OpenStudy (ray10):

yes so now that is \[\large h(x)\] the homogenous solution correct?

OpenStudy (unklerhaukus):

The particular solution will solve \[y_p′′+y_p′+y_p=14+8x+11\cos(x)\] now, it is a good idea to break the particular solution into two parts \[y_p=y_{p_1}+y_{p_2}\] \[y_{p_1}''+y_{p_1}'+y_{p_1}=14+8x\]and\[y_{p_2}''+y_{p_2}'+y_{p_2}=11\cos x\] Now i would recommend undetermined coefficients to find \(y_{p_1}\) and\( y_{p_2}\)

OpenStudy (unklerhaukus):

yeah i think we are using different letters for notation i use \[y(x)=y_c(x)+y_p(x)\] to say the general solution is the sum of the complementary homogenous, and particular solutions but you might have y(x)=h(x)+p(x) or somesuch

OpenStudy (unklerhaukus):

so what would be the 'guess' particular solution (with undetermined coefficients) \[y_{p_1}=h_1(x)=?\] to \[y_{p_1}''+y_{p_1}'+y_{p_1}=14+8x\]

OpenStudy (ray10):

oh of course! yes fair enough so I see how you are looking at it now :) yes I have y(x)=h(x)+p(x)

OpenStudy (unklerhaukus):

ops ** typo \[y_{p_1}=p_1(x)=?\]

OpenStudy (ray10):

ah I was confused there haha I was trying to work it out as well :P

OpenStudy (ray10):

the particular solution guess then?

OpenStudy (unklerhaukus):

Yeah ,

OpenStudy (ray10):

is it \[\large axe^{kx}\] ? I was looking at the trial solution table

OpenStudy (ray10):

this question is really confusing to me :S

OpenStudy (unklerhaukus):

the non-homogenous term in the first particular solution is 14+8x

OpenStudy (ray10):

:O that solution is missing from my table!! :O

OpenStudy (ray10):

and the square box is \[\Large x^{m}\]

OpenStudy (unklerhaukus):

its the easiest case, you can probably remember it , it just says if the non-homogenous term is a polynomial, the guess will be a polynomial of the same order, say the non-homogenous term was 2x^3, the guess would be Ax^3+Bx^2+Cx+D

OpenStudy (ray10):

oh because the power was 3?

OpenStudy (unklerhaukus):

yeah,

OpenStudy (ray10):

and if it was \[\large 2x^{4}\] it woud be \[\large Ax^{4}+Bx^{3}+Cx^{2}+Dx+E\]?

OpenStudy (unklerhaukus):

that's right,

OpenStudy (unklerhaukus):

so what do you get for 14+8x ?

OpenStudy (ray10):

hmmm well its only to the power of one here

OpenStudy (ray10):

correct me if I'm wrong, but does it seem to work in aspect of differentiating?

OpenStudy (ray10):

like following term is m-1

OpenStudy (ray10):

so 14+8x becomes 8?

OpenStudy (unklerhaukus):

there will be constants terms in the 'guess' , we have to solve for these in a moment

OpenStudy (ray10):

that is where I get confused with this question because originally we were working with, getting to \[\large Ae^{kx}+Be^{kx}\]

OpenStudy (unklerhaukus):

?

OpenStudy (ray10):

like what will the guess be? how do we work out the guess for this 14+8x

OpenStudy (unklerhaukus):

our non homogenoues termi is a first order polynomial, So the 'guess' or trial solution well be a first order polynomial with two constants (dont use A, B thought because we already have them in the homogenous solution, use C, D)

OpenStudy (ray10):

14C+8xD, is what I understand from that, have I gone about it wrong? should it be C+xD?

OpenStudy (unklerhaukus):

yes our trial solution for the first particular solution will be \[y_p(x)=p(x)=C+Dx\]

OpenStudy (ray10):

Woooo@@ I'm finally understanding this more :D

OpenStudy (unklerhaukus):

Now take the first and second derivative of p(x)

OpenStudy (ray10):

\[x'= D\]\[x''=0\]

OpenStudy (unklerhaukus):

yeah but they are p' and p''

OpenStudy (ray10):

ohh of course! \[\large p'=D\] \[\large p'=0\]

OpenStudy (unklerhaukus):

actually p'_1, p_''1, *

OpenStudy (unklerhaukus):

ok so now sub them into \[ p_1''+p_1'+p_1=14+8x\]

OpenStudy (ray10):

you mean: \[\large p_{1} ' ,p_{1}''\]? oh you already are a step ahead :D

OpenStudy (ray10):

\[\large p_{1} \] in this case is 14+8x?

OpenStudy (ray10):

\[\large 0+D+14+8x=14+8x\]?

OpenStudy (unklerhaukus):

p_1 was our trial soution p_1=C+Dx

OpenStudy (ray10):

my mistake

OpenStudy (ray10):

\[\large 0+D+C+Dx=14+8x\]

OpenStudy (unklerhaukus):

yes that's better can you solve for C and D?

OpenStudy (ray10):

I get a general solution answer :S

OpenStudy (ray10):

I'll show you

OpenStudy (unklerhaukus):

There is a special technique here you can use , its called equating coefficients \[(C+D)+(D)x=(14)+(8)x\]

OpenStudy (ray10):

oh alright, thank you for showing me that method :)

OpenStudy (unklerhaukus):

if follow that (C+D)=14, D=8

OpenStudy (ray10):

do I get D=8 and C=6?

OpenStudy (unklerhaukus):

yes great!

OpenStudy (ray10):

oh this is great!! :D

OpenStudy (unklerhaukus):

so our trial solution was \[p_1=C+Dx\] we now have C and D so what is the first particular solution equal to?

OpenStudy (ray10):

\[\large p_{1}=C+Dx=6+8x\]

OpenStudy (unklerhaukus):

excellent now all we need is the second particular solution, the one that solves ; \[p_2''+p_2'+p_2=11\cos x\]

OpenStudy (unklerhaukus):

What will our trial solution be this time?

OpenStudy (ray10):

thank you :) Alright so now we must find a particular solution guess correct?

OpenStudy (unklerhaukus):

yeah, you can use a table if you need

OpenStudy (ray10):

\[\large c_{1}\cos(ax+b)+c_{2}\sin(ax+b)\] ?

OpenStudy (unklerhaukus):

that is the right bit of the table to look at, but, our non homogenous term was 11cosx so \[ c_{1}\cos(x)+c_{2}\sin(x)\]or\[ E\cos(x)+F\sin(x)\] is fine

OpenStudy (ray10):

because there is no addition in the cos brackets of course :) yes I understand that :)

OpenStudy (unklerhaukus):

yeah if you really wanted to you could have those extra constants but it would be tonnes more work and the would turn out to be a=1, b=0

OpenStudy (ray10):

I just differentiated them, is that correct?

OpenStudy (unklerhaukus):

Yes, take the first and second derivative of this second particular trial solution , then sub them into \[p_2''+p_2'+p_2=11\cos x\], solve for E, F by equating coefficients ,

OpenStudy (ray10):

for \[\large p_{1}'=E(Fcos(x)-Esin(x))\]\[\large p_{1}''=-E(Ecos(x)+Fsin(x))\]

OpenStudy (unklerhaukus):

wait, what?

OpenStudy (ray10):

then it is \[\large p_{1}''=-E(Ecos(x)+Fsin(x))\] +\[\large p_{1}'=E(Fcos(x)-Esin(x))\]+\[\large p_{1}=E(cos(x)+Fsin(x)\]